OpenStudy (anonymous):

Find the volume of the solid obtained by rotating about x-axis the region bounded by y = cosx and y = 0 and x = 0 and x = pi/2

OpenStudy (anonymous):

\[\int\limits_{0}^{\frac{ \pi }{ 2 }} \pi \cos^2 x dx\]

OpenStudy (anonymous):

∫cos²(x) dx= ∫[1-sin²(x)] dx= =x- ∫sen²(x) dx ∫u dv=u.v-∫vdu u=sin(x) and dv=sin(x) dx, du=cos(x) v=∫sin(x) dx=-cos(x) ∫sin²(x) dx= sin(x) cos(x)+∫cos²(x) dx ∫cos²(x) dx=x-∫sin²(x) dx= =x-sin(x) cos(x)-∫cos²(x) dx 2.∫cos²(x) dx=x-sin(x) cos(x) ∫cos²(x).dx=[x-sin(x).cos(x)]/2 sin(x).cos(x)=sin(2x)/2 ∫cos²(x).dx=x/2-sin(2x)/4 ∫cos²(x).dx=[x-sin(x) cos(x)]/2 [pi/2-sin(pi/2) cos(pi/2)]/2 - [ 0 -sin 0 cos 0]/2 = pi/4 - 0 = pi/4